sam makes $6 per hour at his part time job. so far, he has $120 saved. he is saving all of his earning for computer that costs $498. how many hours dies he need to work in order to buy the computer
Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)



computer costs $498
so far, he has $120 saved.
He needs 498-120 = $378
Sam makes $6 per hour at his part time job.
let number of days be x to earn 378
6x = 378
x = 63







he needs to work for 63 hours in order to buy the computer

Answer by greenestamps(13200)



The answer from the other tutor is undoubtedly what the author of the problem intended.

However, it is not the answer to the question that was asked.

The question that is asked in the problem is "how many hours does he need to work in order to buy the computer?"

The answer is the cost of the computer, divided by how much he earns each hour: 498/6 = 83.

He has already earned $120, which means he has already worked 120/6=20 hours. So he needs to work another 83-20=63 MORE hours to earn enough to buy the computer.

But the problem asks for how many hours he needs to work -- not how many MORE hours he needs to work.